OpenStudy (anonymous):

Initially sliding with a speed of 2.1m/s , a 1.9kg block collides with a spring and compresses it 0.41m before coming to rest. What is the force constant of the spring? answer in units of N/m. Thanks!!

OpenStudy (anonymous):

@Zale101 youve helped me before . can you help me out pleasee!!!!

OpenStudy (surry99):

PE = KE 1/2 kx^2 = 1/2 mv^2 So k = mv^2/x^2 = 1.9 kg (2.1 m/s)^2/ (.41 m)^2 = 49.8 N/m